OpenStudy (anonymous):

6.Factorise the following expressions completely. @ganeshie8

OpenStudy (anonymous):

b) 9+7x-2x^2=

OpenStudy (anonymous):

\( A=-2x^2 + 7x + 9 \) then consider \( -2A = (2x)^2 - 7(2x) - 9(2) \) can u factorise this expression?

OpenStudy (anonymous):

=2x^2-14x-18

OpenStudy (anonymous):

are u sure u have copied the question completely?

OpenStudy (anonymous):

yep @PFEH.1999

OpenStudy (anonymous):

yes it's correct ;)

OpenStudy (anonymous):

forget that way... an awesome way is that find zeroes of that equation : \( A=-2x^2 + 7x + 9 \)

OpenStudy (anonymous):

What u mean by finding zeroes of that equation? @PFEH.1999

OpenStudy (anonymous):

find x in which A=0

OpenStudy (anonymous):

okay..let me think. :)

OpenStudy (anonymous):

u can use quadratic formula.

OpenStudy (anonymous):

i got =(2x-9)(x+1) @PFEH.1999

OpenStudy (anonymous):

but the answer in my book said is (9-2x)(x+1) @PFEH.1999

OpenStudy (anonymous):

(9-2x) or (2x-9) is the same rite? @PFEH.1999

OpenStudy (anonymous):

u know the zeros are \( -1\) and \( \large \frac{9}{2}\) and if u put 9/2 in both equations you will get zero,but consider that in that equation the sign of \( x^2 \) was negative so ur factor should be negative too ;) i didn't notice that by mistake ;)

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

got it :)

OpenStudy (anonymous):

nice ;)

OpenStudy (anonymous):

\(\bf\huge\color{#ff0000}{T}\color{#ff2000}{h}\color{#ff4000}{a}\color{#ff5f00}{n}\color{#ff7f00}{k}~\color{#ffaa00}{Y}\color{#ffd400}{o}\color{#bfff00}{u}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\) @PFEH.1999

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

:)